Five years ago, Chuck was twice as old as Jeanne will be 6 years from now. He is now one year older than 4 times her age. How old was Chuck 5 years ago? How old will Jeanne be 6 years from now?
Answer by stanbon(75887)

Five years ago, Chuck was twice as old as Jeanne will be 6 years from now. He is now one year older than 4 times her age. How old was Chuck 5 years ago? How old will Jeanne be 6 years from now?
---------
Equation(s):
c-5 = 2(j+6)
c = 4j+1
-----
Substitute for "c" and solve for "j"::
4j+1-5 = 2j+12
2j = 16
j = 8 (Jeanne's age now)
Solve for "c":
c = 4j+1
c = 4*8+1 = 33 yrs (Chuck's age now)
-------------------
How old was Chuck 5 years ago?
Ans: 33-5 = 28 yrs
------------------
How old will Jeanne be 6 years from now?
Ans: 8+6 = 14 yrs
===================
Cheers,
Stan H.